# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ 2.3 IVX-Desparsified LASSO
The dependent variable of the auxiliary LASSO regression (2.10) is the standardized instrumental variable (2.8). It is possible to use its original form (2.7). However, the sample s.d. of $\zeta_{j,t}$ passes to infinity if $w_{j,t}$ has a unit root. Thus, without standardizing the instrument, the theoretical order of $\mu$ for a unit root $w_{j,t}$ would be far larger than the order for a stationary regressor. This difference complicates the theoretical justifications. The standardization in (2.8) makes it possible to unify the theoretical convergence rate of $\mu$ regardless of $w_{j,t}$ being a unit root or stationary regressor.
@@
$\widehat{\omega}_{j}^{{\rm XD}}=\dfrac{\sqrt{\widehat{\sigma}_{u}^{2}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}$ (2.12)
where $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}.$ With this point estimator and the associated standard error, we perform the $t$-test for the null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$. Below, we summarize the testing procedure using the XDlasso estimator.
@@
Step1 Obtain $\widehat{\theta}^{{\rm S}}$ from the Slasso regression (2.2). Save the residual $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ and $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. Step2 Obtain the IV $\zeta_{j,t}$ by the transformation (2.7), and standardize it by (2.8). Step3 Run the auxiliary LASSO regression (2.10), and save the residual $\widehat{r}_{j,t}$ in (2.9). Step4 Compute the XDlasso estimator (2.11) and the standard error (2.12). Step5 Obtain the $t$-statistic $t_{j}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{0,j})\big{/}\widehat{\omega}_{j}^{{\rm XD}}.$ (2.13) Reject $\mathbb{H}_{0}$ under the significance level $\alpha$ if $|t_{j}^{{\rm XD}}|>\Phi_{1-\alpha/2}$, where $\Phi_{a}$ is the $100a$-th percentile of the standard normal distribution.
The testing procedure in Algorithm 1 refines the conventional predictive regression inference using modern high dimensional inference techniques. In the following, we elaborate on the necessity of the IVX transformation by explaining the drawback of Dlasso under high dimensional unit roots.
@@ 2.4 Necessity of IVX Transformation
[…]
$\widehat{\psi}^{(j)}$ $=\arg\min_{\psi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{w}_{j,t-1}-W_{-j,t-1}^{\top}\psi)^{2}+\mu\|D_{-j}\psi\|_{1}.$ (2.15)
where the dependent variable in (2.14) is also scaled here to keep in line with (2.9) and (2.10).
When $w_{j,t}$ is a unit root regressor, this standard procedure fails to correct the bias, hindering valid inference. First, (2.15) is a spurious regression as in Granger and Newbold, (1974). In a low dimensional regression where all regressors have unit roots, the standardized least squares counterpart of the regression (2.15) follows a functional central limit theorem (FCLT)
$\widetilde{\psi}^{(j,{\rm OLS})}$ $=$ $D_{-j}\left(\sum_{t=1}^{n}W_{-j,t-1}W_{-j,t-1}^{\top}\right)^{-1}\sum_{t=1}^{n}W_{-j,t-1}\widetilde{w}_{j,t}$ (2.16)
@@
where $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{p})^{\top}$ is a $p$-dimensional Brownian motion, $\sigma_{j}^{*}=\sqrt{\int_{0}^{1}\mathcal{B}_{j}^{2}-(\int_{0}^{1}\mathcal{B}_{j})^{2}}$, and $\mathcal{B}_{-j}=(\mathcal{B}_{k})_{k\neq j}$. In other words, the least square estimator converges in distribution to a nondegenerate random variable. This is the well-known spurious regression phenomenon in unit root regressions. While (2.16) suggests that the limit target coefficients are random and continuously distributed, the asymptotics for the LASSO estimator (2.15) in high dimensions depends on the sparsity of the target coefficients. The randomness of coefficients is contradictory to the sparsity required in LASSO.
Second, as a consequence of the spurious regression, the least squares residual $\widehat{r}_{j,t}^{{\rm OLS}}=\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{(j,{\rm OLS})}$, which under large $n$ is expected to be approximated by $\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{*}$ in view of (2.16), remains highly persistent. Using $\widehat{r}_{j,t}^{{\rm OLS}}$ as the score vector in low dimensions would therefore keep the Stambaugh bias in the noise component ${\rm N}_{j}$. The same issue is also anticipated in the score vector $r_{j,t}$ in (2.14) by LASSO in high dimensions.
@@
In contrast, with the help of the IVX transformation, the score vector in (2.9) achieves (R1) and (R2). Figure 1 provides an illustrative simulation to compare Dlasso with XDlasso. We set $\beta_{1}^{*}$ (the coefficient of the first unit root) as zero, $n=300$, $(p_{x},p_{z})=(150,300)$, and use i.i.d. innovations (See Eq. (4.1)). We generate the data following the DGP (2.1), with a mixture of unit root and stationary regressors as in Section 4.1. Figure 1 displays the histograms of the $t$-statistics over 2000 replications. The density of XDlasso $t$-statistic is well approximated by $\mathcal{N}(0,1)$, whereas the Dlasso $t$-statistic suffers from a substantial bias.
@@ 3 Asymptotic Theory
We now develop the limit theory to shed light on the asymptotic behavior of the XDlasso estimator. The bias correction of XDlasso is mounted on the workhorse estimator Slasso. Slasso’s consistency is a prerequisite for the ensuing maneuver. Unsurprisingly, this paper’s assumptions share similarities with those in Mei and Shi, (2024). We first state our theoretical assumptions and then highlight and explain the differences between the two papers.
Regarding the asymptotic framework, we define the number of regressors $p=p\left(n\right)$ and the sparsity index $s=s\left(n\right)$ as deterministic functions of the sample size $n$. In asymptotic theory, we explicitly send $n\to\infty$, and it is understood that $p(n)\to\infty$ as $n\to\infty$, while $s\left(n\right)$ is allowed to be either fixed or divergent. Recall that the stationary $e_{t}=X_{t}-X_{t-1}\in\mathbb{R}^{p_{x}}$ is the innovation of the unit root regressor. We assume that the stationary high dimensional vector $v_{t}=(e_{t}^{\top},Z_{t}^{\top})^{\top}$ is generated by the linear transformation of innovations $\varepsilon_{t}=(\varepsilon_{k,t})_{k\in[p]}$:
$v_{t}=\Phi\varepsilon_{t}$ (3.1)
@@
$\Pr\left(|\varepsilon_{k,t}|>a\right)$ $\leq C_{\varepsilon}\exp(-a/b_{\varepsilon}),$ (3.3)
for all $k\in[p]$. Furthermore, $\varepsilon_{k,t}$ and $\varepsilon_{\ell,t}$ are independent for all $k\neq\ell.$
@@
and $\mathbb{\mathbb{E}}\left[\sum_{d=-\infty}^{\infty}\varepsilon_{k,t}\varepsilon_{k,t-d}\right]\geq c_{\varepsilon}$ for any $k\in[p].$
The following Assumption 3 depicts the contemporary correlation of $v_{t}$ defined as (3.1). Define $\Omega=\Phi\Phi^{\top}$ where $\Phi$ has appeared in (3.1).
@@
There are absolute constants $c_{\Omega}$, $C_{\Omega}$ and $C_{L}$ such that: (a) $c_{\Omega}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq C_{\Omega}$; (b) $\max_{j\in[p]}\sum_{\ell=1}^{p}\left|\Phi_{j\ell}\right|\leq C_{L}$.
@@
with absolute constants $C_{\zeta}>0$ and $\tau\in(0,1)$. The choice of $\tau$ determines the persistence of the IV $\zeta_{j,t}$, which is elaborated in Remark 5. The following Assumption 4 characterizes the rate of variable dimension $p$ and sparsity index $s$ with respect to the sample size $n$.
@@
In the following, we expound the differences between the assumptions above and those in Mei and Shi, (2024). First, Assumption 1 imposes the m.d.s. and conditional homoskedasticity conditions for the error term $u_{t}$. Three remarks are in order to justify these two conditions.
@@
Though the m.d.s. assumption is not required in Mei and Shi, (2024), in this paper it is essential for the asymptotic normality of XDlasso. Without the m.d.s. assumption, we would call for long-run covariances that not only complicate the procedures but also rule out stationary regressors in the theory. See Phillips and Lee, (2016, Remark 2.3) for detailed discussions.
@@
In empirical finance, the m.d.s assumption is commonly imposed on the error term, especially for testing asset return predictability (Zhu et al.,, 2014; Kostakis et al.,, 2015). It indicates that the dependent variable (financial asset return) is not predictable if the null hypothesis of zero regression coefficients is not rejected, which aligns with the efficient market hypothesis. In macroeconomic applications, the high dimensional covariates with different degrees of persistence alleviate the concern of variable omission, which makes the m.d.s. assumption plausible. [^3] Mild model misspecification with approximate sparsity can be accommodated by our framework; see Belloni et al., (2012) and Mei and Shi, (2024, Remark 1).
@@
It is possible to extend our methodology and theoretical results to conditional heteroskedastic errors. Under low dimensional predictive regression, Kostakis et al., (2015, Theorem 1) show that the homoskedastic-only standard error of the IVX estimator is robust to conditional heteroskedastic error terms if the regressor of interest is persistent. We conjecture that this result applies to XDlasso, and the expression of our standard error (2.12) is robust to conditional heteroskedasticity in our two empirical applications, where each predictor of interest is persistent. A complete theory for conditional heteroskedasticity in high dimensional predictive regression will deserve a standalone paper for future research.
Second, Mei and Shi, (2024) assumed the innovations follow linear processes, and imposed the mixing condition and the lower bounded long run variances through the coefficients in the linear process. In contrast, our Assumption 2 does not assume any specific form of the linear process for $u_{t}$ and $\varepsilon_{t}$, but directly imposes the exponentially decaying rate for the mixing coefficient and the lower bound of the long run variances.
Third, Assumption 4(a) follows Mei and Shi, (2024) that allows the variable dimension $p$ to diverge at a polynomial rate of the sample size $n$; it can be extended to an exponential rate of $n$ at the cost of expositional complications. Assumption 4(b) imposes a more restrictive condition for the sparsity index $s$, compared to $s=o(n^{1/4})$ in Mei and Shi, (2024), which is reasonable as asymptotic normality is more delicate and demanding than consistency. This condition ensures that the shrinkage bias is accurately estimated to achieve the result (R2) in Section 2.2, so that XDlasso asymptotically follows a normal distribution centered at the true coefficient; also see the second term of (3.13) below.
@@
In practice the sparsity index $s$ is unknown. We therefore recommend $\tau=1/2$ for practitioners, under which the quantity $(\tau\wedge(1-\tau))/4$ achieves its maximum $1/8$ and thus permits the weakest sparsity condition. This is different from the conventional wisdom of IVX (Phillips and Lee,, 2016; Kostakis et al.,, 2015) where $\tau$ is recommended to be as large as 0.95 to minimize the loss of rate efficiency (or local power). The context is different in our setting of high dimensional predictive regression, under which An excessively large $\tau$ will require a very small sparsity index $s$ relative to $n$ and thus cause a severe distortion in the size of the test under finite sample. As discussed in Theorem 2 below, the hypothesis testing based on XDlasso is consistent for a wide class of local-to-zero $\theta_{j}^{*}$, despite a slower convergence rate compared with the case when $\tau$ is close to 1.
By Assumptions 14, we first display a lemma for the consistency of the initial workhorse Slasso estimator. For ease of exposition, we present the convergence rate of Slasso for the coefficients standardized by the sample standard deviations.

### v3

@@ 2.3 IVX-Desparsified LASSO
The dependent variable of the auxiliary LASSO regression (2.12) is the standardized instrumental variable (2.10). It is possible to use its original form (2.9). However, the s.d. of $\zeta_{j,t}$ passes to infinity if $w_{j,t}$ is an LUR. Thus, without standardizing the instrument, the theoretical order of the tuning parameter $\mu_{j}$ in (2.12) for an LUR $w_{j,t}$ would be much larger than the order for a stationary regressor. This difference complicates the theoretical justifications. The standardization in (2.10) unifies the convergence rate of $\mu_{j}$ regardless of $w_{j,t}$ being an LUR or a stationary regressor.
Remark 3.
The central idea of debiasing technique via orthogonalization in high dimensional regressions is shared by the van2014asymptotically, javanmard2014confidence and chernozhukov2018. DML by chernozhukov2018 provides a more general framework by allowing nonlinear and semiparametric models. Given cross sectional data, DML advocates using sample splitting to eliminate $c^{*}$ (in the notation of chernozhukov2018), even though empirical process theory implies that $c^{*}$ should vanish asymptotically. In other words, the quality of finite sample approximation using the sample splitting technique is much better than without it. Justifying sample splitting in time series is not as straightforward as that in cross sectional data (beutner2021justification; adamek2022lasso, Remark 4). Without sample splitting, the theory of empirical processes with nonstationary data will be challenging, and this topic deserves thorough investigations in future research.
@@
$\widehat{\omega}_{j}^{{\rm XD}}=\dfrac{\widehat{\sigma}_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}$ (2.14)
where $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. For simplicity, we focus on homoskedastic errors in the main text. In Section B.3 of the appendix, we provide a formula of the heteroskedasticity-robust standard error, and show that our method is robust to conditional heteroskedasticity in both Monte Carlo simulations (Section B.3) and empirical applications (Section C.2).
With the point estimator (2.13) and the associated standard error in (2.14), we perform the $t$-test for the null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$. We summarize the testing procedure using the XDlasso estimator below.
@@
Step1 Obtain $\widehat{\theta}^{{\rm S}}$ from the Slasso regression (2.4). Save the residual $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ and $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. Step2 Obtain the IV $\zeta_{j,t}$ by the transformation (2.9), and standardize it by (2.10). Step3 Run the auxiliary LASSO regression (2.12), and save the residual $\widehat{r}_{j,t}$ in (2.11). Step4 Compute the XDlasso estimator (2.13) and the standard error (2.14). Step5 Obtain the $t$-statistic $t_{j}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{0,j})\big/\widehat{\omega}_{j}^{{\rm XD}}.$ (2.15) Reject $\mathbb{H}_{0}$ under the significance level $\alpha$ if $|t_{j}^{{\rm XD}}|>\Phi_{1-\alpha/2}$, where $\Phi_{1-\alpha/2}$ is the $100(1-\alpha/2)$-th percentile of the standard normal distribution.
The testing procedure in Algorithm 1 refines the conventional predictive regression inference using modern high dimensional inference techniques. In the following, we elaborate on the necessity of the IVX transformation by explaining the drawback of Dlasso under high dimensional LURs.
@@ 2.4 Necessity of IVX Transformation
$r_{j,t}$ $=\widetilde{w}_{j,t}-W_{-j,t}^{\top}\widehat{\psi}^{(j)}\text{ where }\text{$\widetilde{w}_{j,t}=w_{j,t}/\widehat{\sigma}_{j}$,}$ (2.16)
$\widehat{\psi}^{(j)}$ $=\arg\min_{\psi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{w}_{j,t-1}-W_{-j,t-1}^{\top}\psi)^{2}+\mu_{j}\|D_{-j}\psi\|_{1}.$ (2.17)
where the dependent variable in (2.16) is also scaled to be kept in line with (2.11) and (2.12). For simplicity, in this subsection we temporarily restrict all nonstationary regressors to be pure unit roots ($\rho_{j}^{*}=1$). We show that the standard Dlasso procedure fails to correct the bias in this special case of LURs, thereby invalidating the inference in general cases.
First, (2.17) is a spurious regression as in granger1974spurious. In a low dimensional regression where all regressors have unit roots, the standardized least squares counterpart of the regression (2.17) follows a functional central limit theorem (FCLT)
$\widetilde{\psi}^{(j,{\rm OLS})}$ $=$ $D_{-j}\left(\sum_{t=1}^{n}W_{-j,t-1}W_{-j,t-1}^{\top}\right)^{-1}\sum_{t=1}^{n}W_{-j,t-1}\widetilde{w}_{j,t-1}$ (2.18)
@@
where $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{p})^{\top}$ is a $p$-dimensional Brownian motion, $\sigma_{j}^{*}=\sqrt{\int_{0}^{1}\mathcal{B}_{j}^{2}-(\int_{0}^{1}\mathcal{B}_{j})^{2}}$, and $\mathcal{B}_{-j}=(\mathcal{B}_{k})_{k\neq j}$. In other words, the OLS estimator converges in distribution to a nondegenerate random variable. This is the well-known spurious regression phenomenon in unit root regressions. While (2.18) suggests that the limit target coefficients are random and continuously distributed, the asymptotics for the LASSO estimator (2.17) in high dimensions depends on the sparsity of the target coefficients. The randomness of coefficients is contradictory to the sparsity required in LASSO.
Second, the score vector in (2.16) cannot remove the Stambaugh bias. Even in low dimensions, the least squares residual $\widehat{r}_{j,t}^{{\rm OLS}}=\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{(j,{\rm OLS})}$remains highly persistent due to spurious regression. Using $\widehat{r}_{j,t}^{{\rm OLS}}$ as the score vector in low dimensions would therefore keep the Stambaugh bias in the noise component ${\rm N}_{j}$. This issue in low dimension is also present in the score vector $r_{j,t}$ construction in (2.16) by high dimensional LASSO.
@@
In contrast, with the help of the IVX transformation, the score vector in (2.11) achieves (R1) and (R2). Figure 1 provides an illustrative simulation to compare Dlasso with XDlasso. We set $\beta_{1}^{*}$ (the coefficient of the first unit root) as zero, $n=300$, $(p_{x},p_{z})=(150,300)$, and use i.i.d. innovations (See Eq. (4.1)). We generate the data following the DGP (2.1), with a mixture of stationary and nonstationary regressors as in Section 4.1. Figure 1 displays the histograms of the $t$-statistics over 2000 replications. The density of XDlasso $t$-statistic is well approximated by $\mathcal{N}(0,1)$, whereas the Dlasso $t$-statistic suffers from a substantial bias.
2.5 Joint Inference for Low-Dimensional Coefficients
We have devised the XDlasso inference for a scalar coefficient $\theta_{j}^{*}$. In low dimensional predictive regression, the IVX estimator is applicable to multiple coefficients, which jointly follows an asymptotic multivariate normal distribution. Therefore, a Wald statistic by IVX is available to jointly test the predictability of multiple regressors, provided there are a finite number of them. This test statistic is shown to be valid, even when the parameters of interest involve both stationary and nonstationary regressors.
The validity of Wald test extends to XDlasso in high dimensional predictive regression. Specifically, suppose that we are interested in a subset of regressors indexed by $\mathcal{J}\subset[p]$ with a fixed cardinality $|\mathcal{J}|$. In this case, the XDlasso estimators $\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}})_{j\in\mathcal{J}}$ defined in (2.13) asymptotically follow a multivariate normal distribution. Let the null hypothesis $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}}$ involve $|\mathcal{J}|$ restrictions, where $\theta_{0,\mathcal{J}}$ is a $|\mathcal{J}|$-dimensional vector. We construct the following Wald statistic
${\rm Wald}_{\mathcal{J}}^{{\rm XD}}=(\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}-\theta_{0,\mathcal{J}})^{\top}[\widehat{\Omega}_{\mathcal{J}}^{{\rm XD}}]^{-1}(\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}-\theta_{0,\mathcal{J}}),$ (2.19)
where $\widehat{\Omega}_{\mathcal{J}}^{{\rm XD}}$ estimates the covariance matrix of $\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}$, with its $(j,k)$-th entry
$\widehat{\Omega}_{j,k}^{{\rm XD}}=\widehat{\sigma}_{u}^{2}\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}\widehat{r}_{k,t-1}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\cdot\sum_{t=1}^{n}\widehat{r}_{k,t-1}w_{k,t-1}}$
measuring the covariance between $\widehat{\theta}_{j}^{{\rm XD}}$ and $\widehat{\theta}_{k}^{{\rm XD}}$. Under the null hypothesis, the Wald statistic in (2.19) will follow an asymptotic $\chi^{2}$ distribution with the degree of freedom $|\mathcal{J}|$, enabling joint inference on $\theta_{\mathcal{J}}^{*}$.
@@ 3 Asymptotic Theory
This section develops the limit theory to shed light on the asymptotic behaviors of the XDlasso estimator. Unsurprisingly, this paper’s assumptions share similarities with those in MS24. We state our theoretical assumptions and then highlight and explain the differences between the assumptions in these two papers. Regarding the asymptotic framework, we define the number of regressors $p=p\left(n\right)$ and the sparsity index $s=s\left(n\right)$ as deterministic functions of the sample size $n$. In asymptotic statements, we will explicitly send $n\to\infty$, and it is understood that $p(n)\to\infty$ as $n\to\infty$, while $s\left(n\right)$ is allowed to be either fixed or divergent. Recall that the stationary vector $e_{t}\in\mathbb{R}^{p_{x}}$ is the innovation of the LUR regressors. We assume that the stationary high dimensional vector $v_{t}=(e_{t}^{\top},Z_{t}^{\top})^{\top}$ is generated by the innovations $\varepsilon_{t}=(\varepsilon_{k,t})_{k\in[p]}$ via a linear transformation
$v_{t}=\Phi\varepsilon_{t},$ (3.1)
@@
$\Pr\left(|\varepsilon_{k,t}|>a\right)$ $\leq C_{\varepsilon}\exp(-a/b_{\varepsilon}),\ \ \forall k\in[p].$ (3.3)
Furthermore, $\{\varepsilon_{k,t}\}_{t\in\mathbb{Z}}$ and $\{\varepsilon_{\ell,t}\}_{t\in\mathbb{Z}}$ are independent for all $k\neq\ell.$
Remark 4.
In Assumption 1, Eqs. (3.2) and (3.3) impose the sub-exponential tails for the innovations, which includes the familiar sub-Gaussian tail as a special case. The sub-exponential condition is also imposed by MS24 and it is needed in Section 3.1 to deduce the restricted eigenvalue and deviation bound under high dimensional nonstationary data. The heavy-tail features of financial data like extreme returns are not covered in the current paper, and will be an important extension in future studies.
@@
and the long-run variance $\mathbb{\mathbb{E}}\left[\sum_{d=-\infty}^{\infty}\varepsilon_{k,t}\varepsilon_{k,t-d}\right]\geq c_{\varepsilon}$ for all $k\in[p].$
The following Assumption 3 depicts the contemporary correlation of $v_{t}$ defined as (3.1), as well as the constants in the local-to-unity AR coefficients specified in (2.3). Define $\Omega=\Phi\Phi^{\top}$ where $\Phi$ has appeared in (3.1).
@@
There are absolute constants $\underline{c}$ and $\overline{C}$ such that: (a) $\underline{c}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\overline{C}$; (b) $\max_{j\in[p]}\sum_{\ell=1}^{p}\left|\Phi_{j\ell}\right|\leq\overline{C}$; (c) $\max_{j\in[p_{x}]}|c_{j}^{*}|\leq\overline{C}.$
@@
with absolute constants $C_{\zeta}>0$ and $\tau\in(0,1)$. The choice of $\tau$ determines the persistence of the IV $\zeta_{j,t}$, which will be elaborated in Remark 8. The following Assumption 4 characterizes the number of regressors $p$ and the sparsity index $s$ relative to the sample size $n$.
@@
In the following, we expound the differences between the assumptions above and those in MS24. First, Assumption 1 imposes the m.d.s. and conditional homoskedasticity conditions for the error term $u_{t}$. Three remarks are in order to justify these two conditions.
@@
Although the m.d.s. assumption is not required in MS24, in this paper it is essential for the asymptotic normality of XDlasso. Without the m.d.s. assumption, we would need to use long-run covariances that not only complicate the procedures but also rule out stationary regressors in the theory. See phillips2016robust for detailed discussions.
@@
In empirical finance, the m.d.s. assumption is commonly imposed on the error term, especially when testing asset return predictability (zhu2014predictive; kostakis2015robust). It indicates that the dependent variable (financial asset return) is not predictable if the null hypothesis of zero regression coefficients is not rejected, which aligns with the Efficient Market Hypothesis. In macroeconomic applications, the high dimensional covariates with different degrees of persistence alleviate the concern of variable omission, which makes the m.d.s. assumption plausible. [^2] Mild model misspecification with approximate sparsity can be accommodated by our framework; see belloni2012sparse and mei2022lasso.
@@
It is possible to extend our methodology and theoretical results to conditional heteroskedastic errors. Under low dimensional predictive regressions, kostakis2015robust show that the homoskedastic-only standard error of the IVX estimator is robust to conditional heteroskedastic error terms when the regressor of interest is persistent. We conjecture that this result applies to XDlasso, and the expression of our standard error (2.14) is robust to conditional heteroskedasticity in our two empirical applications, where each predictor of interest is persistent. Simulation results in Tables B.5 and B.6 provide supportive evidence on the conjecture. In Appendix B.3, we also consider a heteroskedasticity-robust standard error (B.5), and verify its validity by simulations. A complete theory of conditional heteroskedasticity in high dimensional predictive regression deserves a standalone paper for future research.
Second, MS24 assume the innovations follow linear processes, and impose the mixing condition and the lower bounded long run variances through the coefficients in the linear process. In contrast, our Assumption 2 does not assume any specific form of the linear process for $u_{t}$ and $\varepsilon_{t}$, but directly imposes the exponentially decaying rate for the mixing coefficient and the lower bound of the long run variances.
Third, Assumption 4(a) follows MS24 by allowing $p$ to diverge at a polynomial rate of $n$; it can be extended to an exponential rate of $n$ at the cost of expositional complications. Assumption 4(b) imposes a more restrictive condition for the sparsity index $s$, compared to $s=o(n^{1/4})$ in MS24. This is understandable as asymptotic normality is more delicate and demanding than consistency. This condition ensures that the shrinkage bias is accurately estimated to achieve the result (R2) in Section 2.2, so that XDlasso asymptotically follows a normal distribution centered at the true coefficient; also see the second term of (3.17) below.
@@
In practice, the sparsity index $s$ is unknown. We therefore recommend $\tau=1/2$ for practitioners, under which the quantity $(\tau\wedge(1-\tau))/4$ achieves its maximum $1/8$ and thus permits the weakest sparsity condition. This is different from the conventional wisdom of IVX (phillips2016robust; kostakis2015robust) where $\tau$ is recommended to be as large as 0.95 to minimize the loss of rate efficiency (or local power). Their context is different from our setting of high dimensional predictive regression, under which an excessively large $\tau$ will permit a very small sparsity index $s$ relative to $n$ and thus cause a severe distortion in the size of the test under finite sample. As discussed in Theorem 2 below, the hypothesis testing based on XDlasso is consistent for a wide class of local-to-zero $\theta_{j}^{*}$, despite a slower convergence rate compared with the case when $\tau$ is close to 1 (campbell2006efficient).
Remark 9.
Our theoretical results do not cover cointegrated regressors. Theorem 4 in MS24 shows that in the presence of cointegration, Slasso over-penalizes the coefficients of cointegrated regressors and shrinks them all the way to zero, regardless of their true values. We are unaware of any regularization method that achieves consistent estimation in $p>n$ regime with cointegrated variables mixed with LURs and stationary ones. Without consistency, inferential theory on the cointegrated variable’s true parameter is beyond reach at this moment. Our Appendix B.2 uses a numerical example to show that under some specifications XDlasso can remain robust despite the presence of cointegrated control variables.
Assumptions 14 will be sufficient for the consistency of Slasso with high dimensional LURs and stationary regressors. As the bias correction of XDlasso is mounted on the workhorse estimator Slasso, the consistency of the latter is a prerequisite for the ensuing maneuver.
3.1 Consistency of Slasso
The leading case of persistent regressors is LUR in the low-dimensional predictive regressions (campbell2006efficient). LUR includes the unit root as a special case, and is thus more general in modeling nonstationary behaviors. lee2022lasso study the variable selection properties by the adaptive LASSO under a finite number of LUR regressors, and MS24 cover the consistency of Slasso under high dimensional unit roots. In this paper, the first theoretical result extends Slasso’s consistency in the latter paper to incorporate the LUR processes in the former one. This generalization calls for sophisticated arguments, as briefed in Remark 10.
The consistency of Slasso is founded on two building blocks. The first one, which is essential and challenging, is the restricted eigenvalue (RE) condition of the Gram matrix of the standardized regressors. For any $L>1$, the RE of any $p\times p$ matrix $\Sigma$ is defined as
$\kappa_{H}(\Sigma,L,s):=\inf_{\delta\in\mathcal{R}(L,s)}\dfrac{\delta^{\top}H^{-1}\Sigma H^{-1}\delta}{\delta^{\top}\delta},$ (3.6)
where $\mathcal{R}(L,s)=\{\delta\in\mathbb{R}^{p}\backslash\{0_{p}\}:\|\delta_{\mathcal{M}^{c}}\|_{1}\leq L\|\delta_{\mathcal{M}}\|_{1},\ \text{for all }|\mathcal{M}|\leq s\}.$ The generic matrix $H$ is a placeholder and varies in different contexts. Let $\widehat{\Sigma}=W^{\top}W/n$ be the sample Gram matrix of all regressors. In the context of Slasso, we consider $\Sigma=\widehat{\Sigma}$ and $H=D$ along with the scale standardization in (2.4). The choice of the constant $L$ is related to the procedures of technical proofs and does not impact the rate of convergence. Following the common practice (bickel2009simultaneous), we set $L=3$ as a convenient choice, and simplify the notation as $\widehat{\kappa}_{D}=\kappa_{D}(\widehat{\Sigma},3,s)$. The quantity $\widehat{\kappa}_{D}$ will appear, according to Lemma 1 in MS24, in the denominator of Slasso’s convergence rates. Therefore, a lower bound for $\widehat{\kappa}_{D}$ is essential for the consistency of Slasso.
The second condition for Slasso’s consistency is the deviation bound (DB) of the cross-product between the error term $u_{t}$ in (2.1) and the standardized regressors. The theoretical order of the tuning parameter $\lambda$ must be no smaller than that of $\|n^{-1}\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\|_{\infty}$ to avoid overfitting. On the other hand, an excessively large $\lambda$ causes over shrinkage and damages consistency. A tight upper bound of $\|n^{-1}\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\|_{\infty}$ is therefore indispensable.
Next, we establish the RE and DB conditions for high dimensional mixed roots, and highlight their similarities to and differences from those in MS24. We then leverage them to derive the convergence rates of Slasso.